# Note on supersymmetric mechanics with spin-orbit interaction

### v1

@@ I Introduction
Supersymmetric mechanics was initially introduced as toy models for supersymmetric field theories [1]. However, it was soon realized that these models are highly relevant to the study of numerous problems in conventional quantum mechanics (for a review, see [2] and references therein). Nevertheless, the superfield approach—inspired by supersymmetric field theories—remains the primary tool for studying of supersymmetric mechanics. Over the past two decades, significant efforts have been devoted to constructing models of $\mathcal{N}\geq 4$ supersymmetric mechanics within this framework. In particular, all $\mathcal{N}=8$ were classified in [3]. Much attention has been devoted to superconformal mechanics (see, e.g., [4] and references therein).
@@ II $\mathcal{N}=2$ supersymmetric systems
$\mathcal{H}=\frac{p^{2}_{r}}{2}+\frac{p^{2}_{\varphi}}{2\nu^{2}r^{2}}+\frac{p_{\varphi}\theta^{1}\theta^{2}}{2r^{2}},\qquad Q=\frac{1}{\sqrt{2}}\left(Q_{1}+\imath Q_{2}\right),\qquad Q_{1}=p_{r}\theta^{1}+\frac{p_{\varphi}\theta^{2}}{\nu r},\quad Q_{2}=p_{r}\theta^{2}-\frac{p_{\varphi}\theta^{1}}{\nu r}\;.$ (24)
@@
$\mathcal{J}=\frac{p_{\varphi}}{\nu}+(1-\nu)\theta^{1}\theta^{2},\quad\Pi=\Pi_{1}+\imath\Pi_{2}=\frac{{\rm e}^{-\imath\nu\varphi}}{\sqrt{2}}\left(p_{r}-\imath\frac{p_{\varphi}-\nu\theta^{1}\theta^{2}}{r}\right)$ (25)

### v2

@@ I Introduction
Supersymmetric mechanics were initially introduced as toy models for supersymmetric field theories [1]. However, it was soon realized that these models are highly relevant to the study of numerous problems in conventional quantum mechanics (for a review, see [2] and references therein). Nevertheless, the superfield approach—inspired by supersymmetric field theories—remains the primary tool for studying of supersymmetric mechanics. Over the past two decades, significant efforts have been devoted to constructing models of $\mathcal{N}\geq 4$ supersymmetric mechanics within this framework. In particular, all $\mathcal{N}=8$ were classified in [3]. Much attention has been devoted to superconformal mechanics (see, e.g., [4] and references therein).
@@ II $\mathcal{N}=2$ supersymmetric systems
$\mathcal{H}=\frac{p^{2}_{r}}{2}+\frac{p^{2}_{\varphi}}{2\nu^{2}r^{2}}+\frac{p_{\varphi}\theta^{0}\theta^{1}}{2r^{2}},\qquad Q=\frac{1}{\sqrt{2}}\left(Q_{0}+\imath Q_{1}\right),\qquad Q_{0}=p_{r}\theta^{0}+\frac{p_{\varphi}\theta^{1}}{\nu r},\quad Q_{2}=p_{r}\theta^{1}-\frac{p_{\varphi}\theta^{0}}{\nu r}\;.$ (24)
@@
$\mathcal{J}=\frac{p_{\varphi}}{\nu}+(1-\nu)\theta^{0}\theta^{1},\quad\Pi=\Pi_{0}+\imath\Pi_{1}=\frac{{\rm e}^{-\imath\nu\varphi}}{\sqrt{2}}\left(p_{r}-\imath\frac{p_{\varphi}-\nu\theta^{0}\theta^{1}}{r}\right)$ (25)
@@ Acknowledgements.
Note added. After the publication of this article in Physics of Particles and Nuclei Letters, we noticed that the model given by expressions (42)–(44) represents a particular realization of the $D(1,2|\alpha)$ superconformal mechanics proposed by Anton Galajinsky in [13].